OpenStudy (anonymous):

For the curve defined by the parametric equations x = 10sin(2t), y = 16cos(2t), with 0 ≤ t ≤ π, evaluate dy/dx at the point (5√3, 8)

OpenStudy (anonymous):

Use the chain rule.

OpenStudy (anonymous):

\[ \frac{dy}{dt} =\frac{dy}{dx}\frac{dx}{dt} \]

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = \frac{ \frac{ dy }{ dt } }{ \frac{ dx }{ dt } }\]

OpenStudy (anonymous):

First Find dy/dt..... then dx/dt....and Divide them